Determine the residue of a. 2x+3/z^2-4 b. z-3/z^3+5z^2
1
Expert's answer
2019-09-03T13:13:08-0400

The residue at 0 of (2z+3)/z2-4=2/z+3/z2-4 is 2 since it is equal to the coefficient a-1 in Laurent expansion around 0.

Analogously, the residue of (z-3)/z3+5z2=1/z2-3/z3+5z2 at 0 is 0 since the coefficient "a_{-1}" in the Laurent expansion around 0 is 0.